OpenStudy (anonymous):

please integrate sin^4 x cos^4 x dx

OpenStudy (anonymous):

i used double angle formula

OpenStudy (anonymous):

ok lemme try

OpenStudy (anonymous):

my book uses double angle for both of them if the powers are even.

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

Sorry, I lead you in the wrong direction The better way to do it is by using a combination of\[\sin x \cos x={1\over2}\sin(2x)\]\[\sin^2x={1\over2}(1-\cos(2x))\]and\[\cos^2x={1\over2}(1+\cos(2x))\]

OpenStudy (turingtest):

Note: you will need to use all three formulas above. Well, maybe not need to, but I did.